If a line has the points (4,y) and (16,8), with a slope of -2/3, what is y?
Answer by Alan3354(69443)

If a line has the points (4,y) and (16,8), with a slope of -2/3, what is y?
--------------------
Slope, m = diffy/diffx = (y-8)/(4-16) = -2/3
(y-8)/12 = -2/3
3(y-8) = -24
y-8 = -8
y = 0